# bibCyclic.bib

### v1

@@ 1.1 Enumerative geometry of quiver varieties
The third fundamental curve count, the capped vertex, is the main object of interest for this paper. In some ways, the capped vertex is simpler than the first two curve counts: it is defined in non-localized $K$-theory and is purely classical for quiver varieties with large enough framing, a phenomenon known as “large framing vanishing” and proven in [Ok17]. It was proven by Smirnov in [Sm16] (and earlier in cohomology in [PPRationality]) that the capped vertex is a rational functions of the quantum parameters (typically called Kähler parameters in this setting). On the other hand, computing the capped vertex explicitly in examples is difficult; the localization theorem essentially reduces the problem to computing both the bare vertex and capping operator. Building on [Sm16], the recent work of Smirnov and the first author [AySm] developed explicit formulas for the capped vertex for what is arguably the most important quiver variety: the Hilbert scheme of points in the plane. Their formulas reveal a beautiful relationship with Macdonald polynomials and give an enumerative-geometric interpretation to certain evaluations of them.
@@ 1.2 The capped descendant vertex
Assuming Conjecture 2.7 holds, we have
@@
Let $\lambda\in\mathcal{P}$ such that $\text{core}_{l}(\lambda)=\emptyset$. If Conjecture 2.7 holds, then
@@
Setting $z_{i}=0$, we recover the wreath analog of the usual evaluation formula for Macdonald polynomials, see Conjecture 2.7. Again, this fact is not a coincidence and will be used in the proof.
Let us briefly comment on the assumed Conjecture 2.7. For clarity, we restate the conjecture here.
Conjecture 1.3 (Conjecture 2.7).
Let $10\leq m\leq l-$ 1 and let $\lambda\in\mathcal{P}$ such that $\text{core}_{l}(\lambda)=\emptyset$. Then
$H_{\lambda}|_{p^{(j)}_{k}=\delta_{0,j}-\delta_{m,j}(-u)^{k}}=\prod_{\begin{subarray}{c}\square=(a,b)\in\lambda\\ a-b\equiv i\mod l\end{subarray}}(1+uq^{b-1}t^{a-1})$
We do not know a good reference for this conjecture. No doubt it is well known to experts in the field [josh] and we do not claim originality. As the wreath analogs of usual Macdonald theory continue to develop, it is reasonable to hope for this conjecture to be proven.
Interestingly, it also appears to be the case (and can be verified experimentally using a computer) that nice evaluation formulas exist for arbitrary cyclic quiver varieties with one framing, i.e., for wreath Macdonald polynomials for partitions with arbitrary core. We are unsure if such formulas or even Conjecture 1.3 for $m\neq 0$ admit “quantum deformations” as in Theorem 1.2.
@@ 1.3 Ingredients in the proof
Given that our large framing vanishing result (Theorem 5.1) holds for $\tau_{m}$ for any $m$ and that Conjecture 2.7 likewise holds for any $m$, it is in our view not unreasonable to hope for a nice formula for $\hat{V}^{\left(\tau_{m}\right)}$ for $m\neq 0$. Nevertheless, as we have said above, we do not know how to generalize the $R$-matrix formulas to apply to these cases.
@@ 2.8 Cauchy identities
It is expected by experts that the following evaluation formula holds.
Conjecture 2.7 ([josh]).
@@
We assume this conjecture holds. We have verified this conjecture on a computer for $1\leq l\leq 4$ and $\lambda$ such that $|\lambda|\leq 8$.
Combining the above conjecture with the Cauchy identity gives the following.
@@ 5 Large framing vanishing
5.1 $\mathbb{C}^{\times}_{q}$-fixed quasimaps and localization

### v2

@@ 1.1 Enumerative geometry of quiver varieties
The third fundamental curve count, the capped vertex, is the main object of interest for this paper. In some ways, the capped vertex is simpler than the first two curve counts: it is defined in non-localized $K$-theory and is purely classical for quiver varieties with large enough framing, a phenomenon known as “large framing vanishing” and proven in [Ok17]. It was proven by Smirnov [Sm16] for the Jordan quiver and by Zhu [zhu3] for general quivers (and earlier in cohomology in [PPRationality]) that the capped vertex is a rational functions of the quantum parameters (typically called Kähler parameters in this setting). On the other hand, computing the capped vertex explicitly in examples is difficult; the localization theorem essentially reduces the problem to computing both the bare vertex and capping operator. Building on [Sm16], the recent work of Smirnov and the first author [AySm] developed explicit formulas for the capped vertex for what is arguably the most important quiver variety: the Hilbert scheme of points in the plane. Their formulas reveal a beautiful relationship with Macdonald polynomials and give an enumerative-geometric interpretation to certain evaluations of them.
@@ 1.2 The capped descendant vertex
Then we have
@@
Let $\lambda\in\mathcal{P}$ such that $\text{core}_{l}(\lambda)=\emptyset$. Then
@@
Setting $z_{i}=0$, we recover the evaluation formula for wreath Macdonald polynomials, see Proposition 2.7. When the core is empty, this states
$H_{\lambda}|_{p^{(j)}_{k}=\delta_{0,j}-\delta_{m,j}(-u)^{k}}=\prod_{\begin{subarray}{c}\square=(a,b)\in\lambda\\ a-b\equiv m\mod l\end{subarray}}(1+uq^{b-1}t^{a-1})$
This formula, and its generalization to arbitrary core, was recently proven by Romero and Wen in [WenRomero].
We are unsure if the evaluation formulas for nonempty core or even the empty core evaluation for $m\neq 0$ admit “quantum deformations” as in Theorem 1.2.
@@ 1.3 Ingredients in the proof
Given that our large framing vanishing result (Theorem 5.1) holds for $\tau_{m}$ for any $m$ and that Proposition 2.7 likewise holds for any $m$, it is in our view not unreasonable to hope for a nice formula for $\hat{V}^{\left(\tau_{m}\right)}$ for $m\neq 0$. Nevertheless, as we have said above, we do not know how to generalize the $R$-matrix formulas to apply to these cases.
@@ 2.8 Cauchy identities
The following evaluation formula was recently proven by Wen and Romero.
Proposition 2.7 ([josh]).
@@
Combining the above proposition with the Cauchy identity gives the following.
@@ 5 Large framing vanishing
5.1 Torus fixed quasimaps and localization